OpenStudy (anonymous):

how do i simplify 3x(x-4y+7)

OpenStudy (anonymous):

3x(x−4y+7) =(3x)(x+(−4y)+7) =(3x)(x)+(3x)(−4y)+(3x)(7) =3x2−12xy+21x =3x2−12xy+21x I'm pretty sure, at least...